# Towards Open Respiratory Acoustic Foundation Models: Pretraining and Benchmarking

### v1

@@ 6 Conclusion and Future Research Directions
In this paper, we present OPERA, the first open-source respiratory acoustic foundation model pretraining and benchmarking system. OPERA offers a unique curated dataset pool, a ready-to-use evaluation portal as well as a thorough analysis of performance across architectures and tasks. We discuss the limitations of our work and how it can serve as a foundation for future explorations:
@@
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 1749 0.659 ± 0.001 0.669 ± 0.002 0.680 ± 0.006 0.665 ± 0.001 0.673 ± 0.001 Fine-tune 1749 0.672 ± 0.039 0.691 ± 0.008 0.710 ± 0.003 0.703 ± 0.003 0.715 ± 0.006 Fine-tune 6648 0.723 ± 0.010 0.723 ± 0.009 0.739 ± 0.008 0.733 ± 0.002 0.735 ± 0.005
@@
(2) Investigating scaling law in respiratory acoustic foundation models. Recent research on foundation models has uncovered their emergent abilities, largely arising from scaling up pretraining data and model size [56]. It is also interesting to study the scaling laws in respiratory acoustic foundation models. Our benchmark can help to quantify how increasing a model’s scale and its training data can significantly enhance performance on downstream tasks. Based on the currently 404 hours of respiratory audio, our OPERA-CT (31M parameters) and OPERA-GT (21M) models surpass the lightweight OPERA-CE model (4M). With the rapid accumulation of respiratory audio datasets [69, 14], more evaluation of scaling laws should be conducted in future.
@@
This work was supported by ERC Project 833296 (EAR), EPSRC Project RELOAD, Nokia Bell Labs through a donation, and the Institute for Life Sciences (IfLS) HEIF Research Stimulus Fund. Y. Z. is additionally supported by the Cambridge Trust Scholarship.
@@ A.1.1 Pretraining Data Demographics
Diversity and representativeness of the training data are important for a generalizable model. We examine the demographic distribution of the five datasets used for model pretraining. The bar plots in Figure 4 and Figure 5 illustrate the age and gender distributions across four of these datasets. While the demographic details of HF Lung are not publicly available, the data includes 35 male and 21 female subjects, with an average age of 66.58 (according to the paper [32]). By integrating these diverse datasets in OPERA, we achieve a more representative and unbiased demographic distribution compared to any single data source. This highlights the importance of uniting varied sources for pretraining a foundational model: not only increasing the number of data samples but also ensuring a more comprehensive distribution.
@@ A.1.2 Downstream Task Description
Here we give a detailed description of all 19 tasks formulated in the OPERA benchmark. The tasks are categorized into three types:
@@ A.4 Additional Evaluation Results
Figure 15: Radar plot of reciprocal ranks on two groups of tasks. [Evelyn: radar plot updated]
@@ A.4.1 Another Metric for Lung Function Estimation Tasks
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T13 FVC (Breath) 0.329 ± 0.338 0.298 ± 0.252 0.299 ± 0.245 0.295 ± 0.222 0.304 ± 0.259 0.278 ± 0.261 0.291 ± 0.247 ✓* T14 FEV1 (Breath) 0.353 ± 0.469 0.394 ± 0.444 0.392 ± 0.480 0.396 ± 0.435 0.399 ± 0.449 0.381 ± 0.447 0.392 ± 0.466 T15 FEV1/FVC (Breath) 0.178 ± 0.219 0.167 ± 0.165 0.164 ± 0.163 0.174 ± 0.177 0.161 ± 0.152 0.166 ± 0.149 0.162 ± 0.150 ✓* T16 FVC (Vowel) 0.277 ± 0.238 0.294 ± 0.246 0.280 ± 0.253 0.292 ± 0.247 0.292 ± 0.233 0.264 ± 0.260 0.293 ± 0.255 ✓* T17 FEV1 (Vowel) 0.342 ± 0.363 0.396 ± 0.446 0.417 ± 0.462 0.402 ± 0.409 0.359 ± 0.372 0.398 ± 0.455 0.368 ± 0.440    * T18 FEV1/FVC (Vowel) 0.175 ± 0.183 0.167 ± 0.164 0.167 ± 0.157 0.176 ± 0.170 0.167 ± 0.153 0.171 ± 0.162 0.167 ± 0.158 ✓* T19 Breathing Rate 0.212 ± 0.080 0.205 ± 0.080 0.207 ± 0.086 0.207 ± 0.084 0.207 ± 0.099 0.193 ± 0.065 0.186 ± 0.071 ✓*
@@ A.4.2 Fine-tuning Performance
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 828 0.886 ± 0.017 0.933 ± 0.005 0.855 ± 0.012 0.872 ± 0.011 0.741 ± 0.011 Fine-tune 828 0.984 ± 0.012 0.980 ± 0.007 0.957 ± 0.024 0.808 ± 0.032 0.986 ± 0.006
@@
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 7468 0.649 ± 0.001 0.702 ± 0.001 0.781 ± 0.000 0.769 ± 0.000 0.742 ± 0.001 Fine-tune 7468 0.981 ± 0.002 0.935 ± 0.004 0.994 ± 0.001 0.981 ± 0.002 0.986 ± 0.003

### v2

@@ 6 Conclusion and Future Research Directions
In this paper, we present OPERA, the first open-source respiratory acoustic foundation model pretraining and benchmarking system. OPERA offers a unique curated dataset pool, a ready-to-use evaluation portal as well as a thorough analysis of performance across architectures and tasks. We discuss the limitations of our work and how it can serve as a foundation for future explorations.
Limitations.
While our benchmark is comprehensive, covering 19 tasks, we note that some labels, such as the COVID-19 test results in Task 5, were obtained and reported by participants themselves. As a result, some labels could be less precise than clinically validated data like Task 2. This issue does not affect the pretrained foundation models, as they are trained without these labels. Additionally, OPERA is not intended for clinical use and should not be considered safe for such applications. Care should be taken to prevent potential misuse when utilizing the models.
In addition to the study we have done, OPERA can support a number of future explorations:
@@
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 1749 0.659 $\pm$ 0.001 0.669 $\pm$ 0.002 0.680 $\pm$ 0.006 0.665 $\pm$ 0.001 0.673 $\pm$ 0.001 Fine-tune 1749 0.672 $\pm$ 0.039 0.691 $\pm$ 0.008 0.710 $\pm$ 0.003 0.703 $\pm$ 0.003 0.715 $\pm$ 0.006 Fine-tune 6648 0.723 $\pm$ 0.010 0.723 $\pm$ 0.009 0.739 $\pm$ 0.008 0.733 $\pm$ 0.002 0.735 $\pm$ 0.005
@@
(2) Investigating scaling law in respiratory acoustic foundation models. Recent research on foundation models has uncovered their emergent abilities, largely arising from scaling up pretraining data and model size [57]. It is also interesting to study the scaling laws in respiratory acoustic foundation models. While the OPERA dataset is already extensive, further expansion would be valuable for this purpose. Our benchmark can help to quantify how increasing a model’s scale and its training data can significantly enhance performance on downstream tasks. Based on the currently 404 hours of respiratory audio, our OPERA-CT (31M parameters) and OPERA-GT (21M) models surpass the lightweight OPERA-CE model (4M). With the rapid accumulation of respiratory audio datasets [70, 14], more evaluation of scaling laws should be conducted in future.
@@
This work was supported by ERC Project 833296 (EAR), EPSRC Project RELOAD, Nokia Bell Labs through a donation, Cambridge Trust, and the Institute for Life Sciences (IfLS) HEIF Research Stimulus Fund.
@@ A.1.1 Pretraining Data Demographics
Diversity and representativeness of the training data are important for a generalizable model. We examine the demographic distribution of the five datasets used for model pretraining. The bar plots in Figure 5 and Figure 6 illustrate the age and gender distributions across four of these datasets. While the demographic details of HF Lung are not publicly available, the data includes 35 male and 21 female subjects, with an average age of 66.58 (according to the paper [32]).
Among the five datasets, COVID-19 Sounds and CoughVID were collected globally, while UK COVID-19 and ICBHI were primarily collected in European countries, and HF Lung was collected in Asian regions. Therefore, our curated data presents a comprehensive geo-distribution, covering participants from different ethnic backgrounds and speaking various languages.
Figure 7 summarizes in detail all demographics and medical conditions for the five datasets used for model pre-training. The five datasets used cover a wide range of respiratory medical conditions. COVID-19 Sounds, UK COVID-19, and CoughVID were collected during the pandemic and include some participants who tested positive or negative for COVID-19. Some of the participants had other conditions such as asthma, COPD, pulmonary fibrosis, cancer, etc. The ICBHI and HF Lung datasets include participants who were either healthy or had various respiratory diseases including asthma, COPD, URTI, Pneumonia, etc. Recordings feature both healthy individuals and those with symptoms such as wheeze, crackles, or rhonchi.
By integrating these diverse datasets in OPERA, we achieve a more representative and unbiased demographic distribution compared to any single data source. This highlights the importance of uniting varied sources for pretraining a foundational model: not only increasing the number of data samples but also ensuring a more comprehensive distribution.
Figure 7: Statistics of demographics and medical conditions for datasets used for pretraining.
@@ A.1.2 Downstream Task Description
Figure 8: Statistics of demographics for downstream tasks.
Here we give a detailed description of all 19 tasks formulated in the OPERA benchmark. The demographic statistics are summarized in Figure 8. The tasks are categorized into three types:
@@ A.4 Additional Evaluation Results
Figure 18: Radar plot of reciprocal ranks on two groups of tasks.
@@ A.4.1 Another Metric for Lung Function Estimation Tasks
ID Task Abbr. Opensmile VGGish AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT T13 FVC (Breath) 0.329 $\pm$ 0.338 0.298 $\pm$ 0.252 0.299 $\pm$ 0.245 0.295 $\pm$ 0.222 0.304 $\pm$ 0.259 0.278 $\pm$ 0.261 0.291 $\pm$ 0.247 ✓* T14 FEV1 (Breath) 0.353 $\pm$ 0.469 0.394 $\pm$ 0.444 0.392 $\pm$ 0.480 0.396 $\pm$ 0.435 0.399 $\pm$ 0.449 0.381 $\pm$ 0.447 0.392 $\pm$ 0.466 T15 FEV1/FVC (Breath) 0.178 $\pm$ 0.219 0.167 $\pm$ 0.165 0.164 $\pm$ 0.163 0.174 $\pm$ 0.177 0.161 $\pm$ 0.152 0.166 $\pm$ 0.149 0.162 $\pm$ 0.150 ✓* T16 FVC (Vowel) 0.277 $\pm$ 0.238 0.294 $\pm$ 0.246 0.280 $\pm$ 0.253 0.292 $\pm$ 0.247 0.292 $\pm$ 0.233 0.264 $\pm$ 0.260 0.293 $\pm$ 0.255 ✓* T17 FEV1 (Vowel) 0.342 $\pm$ 0.363 0.396 $\pm$ 0.446 0.417 $\pm$ 0.462 0.402 $\pm$ 0.409 0.359 $\pm$ 0.372 0.398 $\pm$ 0.455 0.368 $\pm$ 0.440    * T18 FEV1/FVC (Vowel) 0.175 $\pm$ 0.183 0.167 $\pm$ 0.164 0.167 $\pm$ 0.157 0.176 $\pm$ 0.170 0.167 $\pm$ 0.153 0.171 $\pm$ 0.162 0.167 $\pm$ 0.158 ✓* T19 Breathing Rate 0.212 $\pm$ 0.080 0.205 $\pm$ 0.080 0.207 $\pm$ 0.086 0.207 $\pm$ 0.084 0.207 $\pm$ 0.099 0.193 $\pm$ 0.065 0.186 $\pm$ 0.071 ✓*
@@ A.4.2 Fine-tuning Performance
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 828 0.886 $\pm$ 0.017 0.933 $\pm$ 0.005 0.855 $\pm$ 0.012 0.872 $\pm$ 0.011 0.741 $\pm$ 0.011 Fine-tune 828 0.984 $\pm$ 0.012 0.980 $\pm$ 0.007 0.957 $\pm$ 0.024 0.808 $\pm$ 0.032 0.986 $\pm$ 0.006
@@
Method # Train AudioMAE CLAP OPERA-CT OPERA-CE OPERA-GT Linear 7468 0.649 $\pm$ 0.001 0.702 $\pm$ 0.001 0.781 $\pm$ 0.000 0.769 $\pm$ 0.000 0.742 $\pm$ 0.001 Fine-tune 7468 0.981 $\pm$ 0.002 0.935 $\pm$ 0.004 0.994 $\pm$ 0.001 0.981 $\pm$ 0.002 0.986 $\pm$ 0.003
A.4.3 Cross-domain Zero-shot Performance
Zero-shot capacity is an particularly interesting trait for foundation models, especially LLM-based models. Though this is uncommon for models trained solely with unlabeled non-textual data, we also explore cross-domain zero-shot performance following [41]. We train a linear probe on source Task A and test it on target Task B, using T6 $\rightarrow$ T9 and T7 $\rightarrow$ T10 as examples, given their similarity (ref. Table 2). Table below shows that OPERA-CT outperforms the baselines.
Table 13: AUROC (higher is better) for cross domain zero-shot performance. Best model highlighted.
Method Opensmile VGGish AudioMAE CLAP OPERA-CT T6 $\rightarrow$ T9 0.534 $\pm$ 0.048 0.537 $\pm$ 0.025 0.472 $\pm$ 0.003 0.457 $\pm$ 0.005 0.600 $\pm$ 0.009 T7 $\rightarrow$ T10 0.682 $\pm$ 0.014 0.588 $\pm$ 0.002 0.692 $\pm$ 0.003 0.722 $\pm$ 0.002 0.823 $\pm$ 0.001
A.4.4 Performance for different model architectures
To investigate whether models trained using OPERA data consistently outperforms models trained with general audio data, comparison using consistent model architectures is also important. We used the same ViT from AudioMAE in OPERA-GT. Similarly, for CNNs, we pretrained VGG (same as VGGish) and CNN14 (as CLAP) using the contrastive objective. While in the main paper we chose to showcase OPERA-CE for its competitive performance and potential in constrained scenarios, we include the results here. The better performance of our models suggests the superiority of our curated respiratory audio data and pretrained models for respiratory health.
Table 14: Average AUROC (higher is better) for the health condition inference tasks.
Model VGG CNN14 ViT
General audio 0.584 0.676 0.627
OPERA data 0.653 0.692 0.674
A.4.5 Significance tests
We conducted significance tests for all tasks and the p values indicating significance is shown in Table 15. Compared to the baselines, our models show a significant improvement in most cases. When compared to the best baseline, OPERA-CT performs better (a higher average of AUROC) on 8 tasks, with 5 of these improvements being statistically significant. Our github repo also provides an easy-to-use significance test function for benchmarking purposes and further use.
Table 15: P-values for significance tests (t-test) for Tasks 1-12. Significant values are highlighted in yellow (p<0.01). The cases where OPERA models outperform the best baseline are underlined.
Dataset ID Best baseline OPERA-CT OPERA-CE OPERA-GT UK COVID-19 T1 VGGish 0.0001 0.0002 0.4230 T2 CLAP 0.0000 0.0022 0.0000 COVID-19 Sounds T3 CLAP 0.0075 0.2155 0.9161 T4 CLAP 0.0558 0.0000 0.0011 CoughVID T5 CLAP 0.0003 0.0003 0.0000 T6 CLAP 0.0000 0.0000 0.0000 ICBHI T7 CLAP 0.0000 0.0000 0.0000 Coswara T8 CLAP 0.1586 0.8547 0.0000 T9 Opensmile 0.0000 0.0000 0.0000 KAUH T10 CLAP 0.0183 0.0003 0.9875 Respiratory@TR T11 CLAP 0.7182 0.0439 0.4200 SSBPR T12 Opensmile 0.0027 0.9944 0.0000